OpenStudy (anonymous):

A person leaves home at 8:00 a.m. and drives to a destination at a rate of 40 mph. The person returns at a rate of 25 mph and arrives at 2:30 p.m. If x represents how far it is to the destination, then which of the following expressions represents the amount of time to make the return trip? 25x 6.5 - x x/25

OpenStudy (anonymous):

A person leaves home at 8:00 a.m. and drives to a destination at a rate of 40 mph. The person returns at a rate of 25 mph and arrives at 2:30 p.m. How far was the destination? 65 miles 100 miles 130 miles

OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? If x represents how many minutes to do the job together, then how much of the job does the slowest person do? x/15 x/10 15/x

OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? 4 min 5 min 6 min

OpenStudy (anonymous):

One person can do a certain job in six days, a second person can do the same job in two days, and a third person can do this job in three days. How many days will they take to do the job together? 1 day 2 days 6 days

OpenStudy (anonymous):

A tank can be filled by one pipe in four hours and by a second pipe in six hours; and when it is full, the tank can be drained by a third pipe in three hours. If the tank is empty and all three pipes are open, in how many hours will the tank be filled? 8 hours 12 hours 24 hours

OpenStudy (anonymous):

A mixture contains forty ounces of glycol and water, and it is ten percent glycol. The mixture is to be strengthened to twenty-five percent by adding glycol. How much glycol is in the original mixture? 4 oz 10 oz 40 oz

OpenStudy (anonymous):

A mixture contains forty ounces of glycol and water and is ten percent glycol. If the mixture is to be strengthened to twenty-five percent, how much glycol is to be added? 4 oz 8 oz 10 oz

OpenStudy (anonymous):

An automobile's radiator has a capacity of fifteen quarts, and it currently contains twelve quarts of a thirty percent antifreeze solution. How many quarts of pure antifreeze must be added to strengthen the solution to forty percent? 2 quarts 3 quarts 4 quarts

OpenStudy (anonymous):

These should each be their own topic. Ask them one at a time.

OpenStudy (valpey):

Are you just transcribing all of your homework?

OpenStudy (ganpat):

For average speed = Speed = (2xy)/( x + y) x = speed up and y = speed down.. So, Average speed = (2 * 40 * 25)/ (40 + 25) = 30.77... and using formula, Speed = distance/ time 30.77 = x / 6.5 hrs... so x = 200 miles... so this is up and down distance... one side is 100 miles.. Hope that answers all the questions...

OpenStudy (ganpat):

Time for return trip.. x/25..

OpenStudy (nottim):

augh. too many questions. suggestion: do what the guy before said; ask each question individually. Better yet, attempt all of these questions yourself, and ask only the questions you actually have difficulty on.

OpenStudy (anonymous):

:S